The temp of 1mol of gas is increased by1°c at constant pressure .the work done is

Grade:11

1 Answers

Vikas TU
7 years ago
At constant Pressure,
W = nCp(delta T)
C p  = (F/2 + 1)R = > (3/2+1)R = > 5R/2
 
W = 1* 5R * ½ =  > 5R/2 Joules.